Jane earns $41 more a week than twice as much as Bill does. Together they earn $140. How much does each earn?
Answer by jb007mi6(53)

if J is janes salary and B is bills salary then
J=2B+41 (Jane earns $41 more a week than twice as much as Bill does.)
(Together they earn $140.)
this means
B+(2B+41)=140
3B+41=140
3B=140-41
3B=99
B=33
since b+j=140 and b=33
33+j=140
j=107
check your answer to see if it's right